the ancient greeks thought that the most pleasing shape for a rectangle was one for which the ratio of the lengthto the width was approximatly 8 to 5, the golden ratio. the length of a rectangular painting is 2 ft longer than its width, then for what dimentions would the length and width have the golden rule?
Answer by checkley77(12844)

8/5=(X+2)/X
5(X+2)=8X
5X+10=8X
5X-8X=-10
-3X=-10
X=-10/-3
X=3.33 FT. FOR THE WIDTH.
3.33+2=5.33 FT. FOR THE LENGTH.
PROOF:
8/5=5.33/3.33
8*3.33=5*5.33
26.64~26.65